# Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

### v2

@@ Introduction
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [1]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information of material science from the vast ocean of literature remains demanding. To obtain clearer sense of materials properties some structured database project such as Materials Project [2], OQMD [3], and NOMAD [4] were developed. However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [5]. While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations. Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
@@
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [6], [7]. These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [8]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [9]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [10]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [11]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [12]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [13]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [14]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [1], [15].
@@
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first accurate knowledge graph dedicated for materials, where researchers can easily get the information about material through query the MKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs. 4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
@@ Methods
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [15], highlighting the differences in their construction methodologies as illustrated in Figure 4. MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
@@
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts. This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets. Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text. This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis confirms the effectiveness and credibility of the MKG. Furthermore, we have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding. This analysis has enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions. First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG. This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature. Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG. Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science. Lastly, the integration of our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.

### v3

@@ Introduction
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions [1]. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [2]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the availability of comprehensive databases such as Scopus, Web of Science, and Crossref that offer structured search capabilities for scientific literature, extracting actionable materials science information from the extensive volumes of published research remains a daunting task. In response, databases like the Materials Project [3], OQMD [4], and NOMAD [5] have been developed to provide a clearer understanding of materials properties. While traditional databases are effective in storing structured data and handling basic queries, they may struggle to capture complex relations and infer new knowledge from the data [6]. Specifically, traditional databases store data in tables with a predefined schema, organizing data into rows and columns, and establish relations between entities using primary and secondary keys. This architecture often requires complex and computationally intensive queries to extract interconnected data, which can hinder the effective mining of knowledge from various scientific findings.
@@
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability and visualization [7], the ability to infer new knowledge through relational data analysis or graph neural network [8], and improved data quality and consistency through structured representation [9], [10]. These features make knowledge graphs particularly valuable for integrating diverse information sources, providing a unified view of a domain’s knowledge and more comprehensive contextual relationships, thereby facilitating more informed decision-making and discovery [11]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [12]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [13]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [14]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [15]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [16]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [17]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [2], [18].
@@
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are as follows: (1) We propose a method to achieve high accuracy in Named Entity Recognition (NER), Relation Extraction (RE), and Entity Resolution (ER). This method allows for efficient conversion of unstructured text from complex research papers into structured information, effectively capturing crucial information and accelerating the research process in the early stages. This process not only retains the source information for each triplet but also makes updating the KG very convenient. (2) We constructed a dynamic MKG that supports rapid and automated updates, which is invaluable for the aggregation and processing of the growing global data on materials science. This capability ensures that researchers have access to the most current data, enhancing the predictive accuracy as the database expands. (3) We use a well-defined label system that enables easy scaling of our KG and potential integration with other structured databases or KGs like MatKG series and BoschKS [25]. This well-defined labeling not only integrates professionalism into the construction of the knowledge graph but also acts as an effective link between researchers from both academia and industry. (4) We demonstrate a similarity calculation method based on the Jaccard Similarity for materials and applications. This protocol could enable the integration of knowledge and the discovery of novel materials for material applications from the wealth of scientific literature, which can facilitate the design process and inspire innovative ideas for future studies of material science researchers.
@@ Methods
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG series [2], [18], highlighting the differences in their construction methodologies as illustrated in Figure 4. However, the relations in MatKG is based on statistical means, which can not guarantee high accuracy of the relations, especially when the article contains multiple materials. Therefore, our main comparison object is MatKG2. MatKG2 is constructed using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
@@
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts. This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets. Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text. This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted. Moreover, this method can automatically extract high-precision knowledge form corpus, leading to dynamic KGs.
Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis confirms the effectiveness and credibility of the MKG. We have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding, which enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions. First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG. This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature. Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG. Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science. Lastly, the integration of our Material Knowledge Graph with existing general material KGs, such as MatKG series and BoschKS, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.